# A Mixture-of-Experts Framework with Log-Logistic Components for Survival Analysis on Histopathology Images

### v2

@@ Abstract
We propose a modular framework for predicting cancer-specific survival from whole-slide pathology images (WSIs). The method integrates four components:(i) Quantile-Gated Patch Selection via quantile-based thresholding to isolate prognostically informative tissue regions; (ii) Graph-Guided Clustering using $k$-NN graph to capture phenotype-level heterogeneity through spatial–morphological coherence; (iii) Hierarchical Context Attention to learn intra- and inter-cluster interactions; and (iv) an Expert-Driven Mixture of Log-logistics framework to estimate complex survival distributions using Log-logistics distributions. The model attains a concordance index of $0.644{\pm}0.059$ on TCGA-LUAD, $0.751{\pm}0.037$ on TCGA-KIRC, and $0.752{\pm}0.011$ on TCGA-BRCA, respectively, outperforming existing state-of-the-art approaches.
Index Terms—  Survival Analysis, Whole-Slide Images, Histopatholgy, Attention, Mixture of Experts, Deep Learning.
@@ 1 Introduction
Accurate survival prediction from histopathology images is critical for personalized cancer care, guiding treatment and risk stratification. Whole-slide images (WSIs) capture rich tissue-level morphology—including tumor architecture and microenvironmental context—but their massive size and lack of localized annotations make direct modeling challenging. Patch-based multiple instance learning (MIL) has thus become the dominant paradigm, dividing each slide into smaller tiles for weakly supervised analysis.
While MIL has succeeded in cancer grading and subtyping, survival prediction requires modeling long-range dependencies and subtle morphological cues across spatially distributed tissue regions. Conventional statistical approaches such as the Cox proportional hazards model cannot capture these non-linear, high-dimensional relationships, whereas deep models often lack interpretability and spatial reasoning.
To address these challenges, we present a unified framework combining four modules: (1) Quantile-Gated Patch Selection dynamically identifies patches and retains prognostically relevant tissue regions through quantile-based filtering. (2) Graph-Guided Clustering of Relevant Patches to capture spatial morphological coherence; (3) Hierarchical Context Attention to learn both intra- and inter-cluster dependencies; and (4) expert-driven mixture of Log-logistics framework for survival distribution estimation. Together, these components enable interpretable and accurate survival prediction by jointly leveraging spatial structure, phenotype abstraction, and probabilistic outcome modeling.
@@ 2 Related Work
Weakly supervised learning has been central to WSI-based survival modeling, primarily through multiple instance learning (MIL). Early models such as ABMIL [ilse2018attentionbaseddeepmultipleinstance], CLAM [lu2020dataefficientweaklysupervised], and DSMIL [li2021dualstreammultipleinstancelearning] used attention pooling to highlight prognostic patches but overlooked spatial context. Graph-based PatchGCN [chen2021slideimages2dpoint] captured spatial adjacency, while hierarchical transformers like HGT [10.1007/978-3-031-43987-2_72], TransMIL [shao2021transmiltransformerbasedcorrelated], and HIPT [chen2022scalingvisiontransformersgigapixel] modeled multi-scale features at high computational cost. Recent multimodal methods such as PathoGen-X [10981028], CoC [ZhoHai_CoC_MICCAI2025], and HiLa [CuiJia_HiLa_MICCAI2025] integrate histopathology with genomic, methylation, and language data, while SCMIL [Yang_2024] estimates Gaussian distribution-based survival predictions. However, these approaches lack mechanisms to disentangle latent phenotypes or exploit structural coherence. Our expert-driven mixture of log-logistics module addresses this gap by integrating phenotype-aware clustering with expert-specific survival estimation, where specialized experts capture distinct features and a gating network adaptively fuses their outputs for improved interpretability.
Fig. 1: (a) Proposed survival framework. (b) Expert-guided Mixture Density Modeling block.
@@ 3 Methodology
$\mathbf{P}_{\text{feat}}\in\mathrm{R}^{n\times d},$ (1)
where $n$ denotes the number of extracted patches and $d$ the feature dimensionality.
The proposed framework comprises four modules: Quantile-Gated Patch Selection(QGPS), Graph-Guided Clustering(GGC), Hierarchical Context Attention(HCA), and Expert-Driven Mixture of Log-logistics Module(EDMLL).
@@ 3.1 Quantile-Gated Patch Selection (QGPS)
To focus on diagnostically informative regions, a quantile-based learnable Multi Layer Perceptron(MLP) selects task-relevant patches according to their importance scores derived from model logits. For each WSI, a threshold $\tau_{q}$ is determined as the $q$-quantile of logits, retaining the top $(1-q)\times 100\%$ of patches:
@@
This adaptive selection removes noisy or background patches, ensuring downstream modules focus on semantically meaningful regions relevant to survival outcomes.
@@ 3.2 Graph-Guided Clustering (GGC)
The selected patches $\mathcal{P}_{\text{sel}}\in\mathrm{R}^{m\times d}$ are clustered into morphologically and spatially coherent groups. First, patch features and coordinates are normalized to comparable scales. Morphological similarity between patches is computed using cosine similarity:
@@
while spatial similarity is defined via an exponential kernel over Euclidean distances $D_{ij}$ between normalized coordinates:
$S_{\text{spatial}}(i,j)=\exp\!\left(-\frac{D_{ij}}{\sigma_{D}}\right),\quad\text{where }\sigma_{D}=\text{std}(D)+\varepsilon.$ (4)
A composite similarity matrix balances both cues:
$S=\omega_{\text{morph}}S_{\text{morph}}+\omega_{\text{spatial}}S_{\text{spatial}}$ (5)
Based on $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and patches are then clustered into $G$ groups $L_{1},\dots,L_{G}$} using GPU-accelerated K-Means, guided by the graph structure. Each cluster represents a local tissue microenvironment such as tumor epithelium, stroma, or necrosis, promoting a biologically meaningful and spatially coherent organization across the slide.
@@ 3.3 Hierarchical Context Attention (HCA)
To capture relationships within clutsers, a two-level attention mechanism is adopted.
Intra-cluster attention. Each cluster $L_{i}$ undergoes multi-head self-attention (MHSA) [vaswani2017attention] to model fine-grained local dependencies among patches:
@@
This step enhances contextual awareness within morphologically similar regions.
Inter-cluster attention. The refined clusters are summarized into representative embeddings
$R_{i}=\frac{1}{|L^{\prime}_{i}|}\sum_{\mathbf{x}\in L^{\prime}_{i}}\mathbf{x},$ (7)
which are then processed by another MHSA layer to capture global relationships across regions:
@@
Intra-cluster representations are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{g})_{g=1}^{G}$. A broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, derived by averaging $R^{\prime}$ across clusters and expanding to match $\widetilde{P}$, is added residually to form $\widehat{P}$. Finally, $\widehat{P}$ is concatenated with remaining irrelevant patch features to yield $\mathcal{P}_{\text{final}}$.
WSI-level aggregation. Slide-level representation is derived via attention pooling as in AMIL [ilse2018attentionbaseddeepmultipleinstance]
$\mathbf{z}_{\text{WSI}}$ $=\sum_{i}\alpha_{i}\,\mathcal{P}_{\text{final},i},$ (9)
$\alpha_{i}$ $=\text{softmax}\!\big(W_{a}\,\tanh(W_{h}\,\mathcal{P}_{\text{final},i}^{\top})\big),$
where $\mathbf{z}_{\text{WSI}}$ represents the integrated prognostic feature for the WSI.
3.4 Expert-Driven Mixture of Log-logistics
An Expert-Driven Mixture of Log-Logistics module (Figure 1b) models survival probabilities, with each expert specializing in distinct WSI features $\mathbf{z}_{\text{WSI}}$. Built on the mixture-of-experts framework [bishop2006pattern], it uses multiple experts to form Log-logistics Distributions(LLD) that capture heterogeneous survival patterns. The log-logistic probability density is given by
$\mathrm{LLD}(t\mid\alpha,\beta)=\frac{(\beta/\alpha)(t/\alpha)^{\beta-1}}{\big[1+(t/\alpha)^{\beta}\big]^{2}},\quad t>0,$ (10)
where $\alpha,\beta>0$ denote the scale and shape parameters.
The survival time $t$ of a whole slide image is directly modeled on its features $\mathbf{z}_{\text{WSI}}$ derived via the attention pooling layer, as shown in Fig. 1(b). Each expert $e$ in the mixture models the conditional distribution over $t$ using a log-logistic mixture model (LLD):
@@ 3.4 Expert-Driven Mixture of Log-logistics
where each $k$ component follows a log-logistic distribution with scale $\alpha^{(e)}_{k}$ and shape $\beta^{(e)}_{k}$, while mixture weights $\lambda^{(e)}_{k}$ are computed using a softmax over the $K$ components. Learnable global anchors $(P_{\alpha},P_{\beta})\in\mathrm{R}^{K}$, corresponding to scale and shape parameters, are projected for each expert as below:
$\alpha^{(e)}=\text{softplus}\!\big(W^{(e)}_{\alpha}P_{\alpha}\big),\quad\beta^{(e)}=\text{softplus}\!\big(W^{(e)}_{\beta}P_{\beta}\big)$ (12)
A gating network $G(\mathbf{z}_{\text{WSI}})$ assigns softmax weights to experts, producing the overall time probability density function (TPDF) and survival probability function (SPF) as:
@@
3.5 Training Objective and Regularization
The model is trained using a negative log-likelihood loss ($\mathcal{L}_{\text{NLL}}$) that handles both censored ($c=0$ i.e. event not yet occurred) and uncensored ($c=1$ i.e. event occurred) data. For an observed time $t_{d}$ and censoring indicator $c\in\{0,1\}$,
$\mathcal{L}_{\text{NLL}}=-\,c\,\log\!\big(\text{TPDF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)-(1-c)\,\log\!\big(\text{SPF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)$
To promote regularization, a gating entropy loss is introduced($\mathcal{L}_{\text{ent}}$).
$\mathcal{L}_{\text{ent}}=-{\textstyle\sum_{e}}G_{e}(\mathbf{z}_{\text{WSI}})\log\!\big(G_{e}(\mathbf{z}_{\text{WSI}})\big).$ (16)
The overall loss combines the two above terms:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{NLL}}+\lambda_{\text{ent}}\,\mathcal{L}_{\text{ent}}.$ (17)
@@ 4.1 Datasets
We evaluate the framework on three TCGA cohorts: Lung Adenocarcinoma (LUAD; 459 WSIs), Kidney Renal Clear Cell Carcinoma (KIRC; 509 WSIs), and Breast Invasive Carcinoma (BRCA; 956 WSIs) [tcga2013pan]. Each dataset contains WSIs annotated with overall survival time, measured in months from diagnosis to death (uncensored data, c = 1) or last follow-up (censored data, c = 0).
4.2 Training Configuration
@@
The quantile hyperparameter is set to $q=0.25$, retaining the top $75\%$ of patches based on logits. The $k$-NN graph constructed from the similarity matrix $S$ connects each point to its 10 nearest neighbors. Based on the graph connectivity, the patch indices are clustered into $G$ groups of 64 each, and these grouped patches are subsequently processed by an 8-head MHSA within the HCA module. The Expert-Driven Mixture of Log-Logistics module employs five experts, each modeling a log-logistic distribution (LLD) with $K=100$ components. Model is trained for 20 epochs with Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) and a batch size of 1. Results are reported as mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric over 5-fold cross-validation. All experiments are conducted on an NVIDIA A6000 GPU.
5 Results and Analysis
5.1 Quantitative Results
@@
Table 1(a) summarizes performance across TCGA datasets. Our framework consistently achieves the highest mean TDC, indicating superior discrimination and calibration. Histology-based multiple instance learning (MIL) models—AMIL [ilse2018attentionbaseddeepmultipleinstance], CLAM [lu2020dataefficientweaklysupervised], DSMIL [li2021dualstreammultipleinstancelearning], and TransMIL [shao2021transmiltransformerbasedcorrelated] served as backbones for the expert-driven mixture of log-logistics block, with performances reported for comparison. Survival-specific MIL methods such as HIPT [chen2022scalingvisiontransformersgigapixel], Patch-GCN [10981028], HGT [10.1007/978-3-031-43987-2_72], SCMIL [Yang_2024] and OTSurv [RenQin_OTSurv_MICCAI2025] act as baselines. Since multimodal frameworks [10981028, ZhoHai_CoC_MICCAI2025, CuiJia_HiLa_MICCAI2025] lack public code, their results are taken from the respective papers. By combining quantile-based patch selection, graph-guided clustering, hierarchical attention, and expert-driven mixture of log-logistics framework—each expert capturing distinct morphological or contextual cues—our model effectively identifies survival-relevant regions and achieves state-of-the-art performance on all datasets. Remarkably, it even surpasses multimodal systems while relying solely on histopathology images.
5.2 Ablation Studies and Interpretability

### v3

# Spatially-Aware Mixture of Experts with Log-Logistic Survival Modeling for Whole-Slide Images

@@ Abstract
Accurate survival prediction from histopathology whole-slide images (WSIs) remains challenging due to gigapixel resolutions, spatial heterogeneity, and complex survival distributions. We introduce a comprehensive computational pathology framework that addresses these limitations through four synergistic innovations: (1) Quantile-Gated Patch Selection to dynamically identify prognostically relevant regions; (2) Graph-Guided Clustering that groups patches by spatial-morphological similarity to capture phenotypic diversity; (3) Hierarchical Context Attention to model both local tissue interactions and global slide-level context; and (4) an Expert-Driven Mixture of Log-Logistics module that flexibly models complex survival distributions. On large-scale TCGA cohorts, our method achieves state-of-the-art performance, with time-dependent concordance indices of $0.644{\pm}0.059$ on LUAD, $0.751{\pm}0.037$ on KIRC, and $0.752{\pm}0.011$ on BRCA—significantly outperforming both histology-only and multimodal benchmarks. The framework provides improved calibration and interpretability, advancing the potential of WSIs for personalized cancer prognosis.
Index Terms—  Survival Analysis, Whole-Slide Images, Histopatholgy, Attention, Mixture of Experts.
@@ 1 Introduction
Accurate survival prediction from histopathology images is critical for personalized cancer care, guiding treatment selection and risk stratification. While clinical and genomic markers are routinely used, histopathology images contain rich, underutilized morphological information that could significantly enhance prognostic accuracy. Whole-slide images (WSIs) capture comprehensive tissue-level morphology—including tumor architecture, stromal composition, and microenvironmental context—but their gigapixel resolutions and lack of localized annotations make direct modeling computationally prohibitive. Patch-based multiple instance learning (MIL) has thus emerged as the dominant paradigm for weakly supervised analysis of WSIs.
However, survival prediction introduces unique challenges beyond conventional classification tasks. It requires modeling long-range spatial dependencies and subtle morphological patterns across distributed tissue regions, where the arrangement and interaction of different tissue phenotypes may carry critical prognostic significance. Conventional statistical approaches, such as Cox proportional hazards, struggle to capture the complex, high-dimensional, non-linear relationships in histopathology data, while deep learning methods often lack interpretability and fail to maintain spatial coherence across the slide.
Current MIL approaches for survival analysis face several limitations: they often treat patches as independent instances, ignoring spatial relationships; they lack mechanisms to disentangle heterogeneous tissue phenotypes; and they typically rely on simplistic survival modeling that cannot capture complex, multi-modal hazard distributions. These limitations restrict their ability to fully leverage the rich morphological information embedded in WSIs.
To overcome these challenges, we introduce a unified computational pathology framework that integrates four key innovations: (1) Quantile-Gated Patch Selection to dynamically identify prognostically relevant regions while filtering noise; (2) Graph-Guided Clustering to group patches by spatial-morphological similarity, capturing phenotypic diversity; (3) Hierarchical Context Attention to model both fine-grained local interactions and global slide-level context; and (4) an Expert-Driven Mixture of Log-Logistics that flexibly models complex survival distributions through specialized components. Together, these elements enable interpretable and accurate survival prediction by jointly leveraging spatial structure, phenotypic abstraction, and probabilistic outcome modeling, providing a more comprehensive approach to WSI-based prognosis.
@@ 2 Related Work
Weakly supervised learning via multiple instance learning (MIL) has been central to WSI-based survival modeling, addressing the challenge of slide-level labels with patch-level features. Early attention-based methods, such as ABMIL [ilse2018attentionbaseddeepmultipleinstance], CLAM [lu2020dataefficientweaklysupervised], and DSMIL [li2021dualstreammultipleinstancelearning], identified prognostic patches but overlooked spatial context and inter-patch relationships. Subsequent approaches incorporated structural information through graph-based methods (PatchGCN [chen2021slideimages2dpoint]) and hierarchical transformers (HGT [10.1007/978-3-031-43987-2_72], TransMIL [shao2021transmiltransformerbasedcorrelated], HIPT [chen2022scalingvisiontransformersgigapixel]), capturing spatial and multi-scale relations at higher computational cost.
Recent efforts have explored multimodal integration of histopathology with genomic (PathoGen-X [10981028]), methylation (CoC [ZhoHai_CoC_MICCAI2025]), and language data (HiLa [CuiJia_HiLa_MICCAI2025]). In parallel, probabilistic methods, such as SCMIL [Yang_2024] employ Gaussian distributions for survival estimation but lack flexibility for complex hazard shapes. However, these approaches typically lack explicit mechanisms for phenotype disentanglement and struggle to model heterogeneous compositions in cancer microenvironments.
Our method addresses these limitations through integrated phenotype-aware clustering and expert-driven survival modeling. We combine graph-guided clustering for spatial-morphological coherence, hierarchical attention for multi-scale context, and a mixture of log-logistic experts that adapt to diverse survival patterns. This unified approach enables interpretable phenotype disentanglement while flexibly modeling complex survival distributions beyond conventional parametric assumptions.
Fig. 1: (a) Proposed survival framework. (b) Expert-Driven Mixture of Log-logistics block.
@@ 3.1 Quantile-Gated Patch Selection (QGPS)
A quantile-based MLP selects task-relevant patches using importance scores from model logits, filtering out uninformative regions to reduce computational burden and noise. For each WSI, threshold $\tau_{q}$ is the $q$-quantile of logits, retaining the top $(1-q)\times 100\%$ of patches:
@@ 3.2 Graph-Guided Clustering (GGC)
Selected patches $\mathcal{P}_{\text{sel}}\in\mathrm{R}^{m\times d}$ are clustered into morphologically and spatially coherent groups to capture phenotypic diversity. Morphological similarity uses cosine similarity:
@@
while spatial similarity uses an exponential kernel over Euclidean distances:
$S_{\text{spatial}}(i,j)=\exp\!\left(-\frac{D_{ij}}{\sigma_{D}}\right),\quad\sigma_{D}=\text{std}(D)+\varepsilon.$ (3)
The composite similarity $S=\omega_{\text{morph}}S_{\text{morph}}+\omega_{\text{spatial}}S_{\text{spatial}}$ guides a $k$-NN graph construction, with patches clustered into $G$ groups $L_{1},\dots,L_{G}$} using GPU-accelerated K-Means.
@@ 3.3 Hierarchical Context Attention (HCA)
Each cluster $L_{i}$ undergoes multi-head self-attention (MHSA) [vaswani2017attention] for local dependencies, capturing fine-grained relationships within phenotypically similar regions:
@@
Refined clusters are summarized into embeddings $R_{i}=\frac{1}{|L^{\prime}_{i}|}\sum_{\mathbf{x}\in L^{\prime}_{i}}\mathbf{x}$, then processed by another MHSA layer for global relationships across different tissue regions:
@@
Intra-cluster representations are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{g})_{g=1}^{G}$. A broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, derived by averaging $R^{\prime}$ across clusters, is added residually to form $\widehat{P}$, which is concatenated with remaining patches to yield $\mathcal{P}_{\text{final}}$.
Slide-level representation is derived via attention pooling:
$\mathbf{z}_{\text{WSI}}=\sum_{i}\alpha_{i}\,\mathcal{P}_{\text{final},i},\quad\alpha_{i}=\text{softmax}\!\big(W_{a}\,\tanh(W_{h}\,\mathcal{P}_{\text{final},i}^{\top})\big),$ (6)
where $\mathbf{z}_{\text{WSI}}$ represents the integrated prognostic feature for the entire slide.
3.4 Expert-Driven Mixture of Log-logistics (EDMLL)
The module models survival probabilities using multiple experts specializing in distinct WSI features $\mathbf{z}_{\text{WSI}}$, enabling capture of diverse risk patterns. Each expert models the conditional distribution using log-logistic mixtures:
@@ 3.4 Expert-Driven Mixture of Log-logistics (EDMLL)
where the log-logistic density is:
$\mathrm{LLD}(t\mid\alpha,\beta)=\frac{(\beta/\alpha)(t/\alpha)^{\beta-1}}{\big[1+(t/\alpha)^{\beta}\big]^{2}},\quad t>0.$ (8)
Learnable global anchors $(P_{\alpha},P_{\beta})\in\mathrm{R}^{K}$ are projected as:
$\alpha^{(e)}=\text{softplus}\!\big(W^{(e)}_{\alpha}P_{\alpha}\big),\quad\beta^{(e)}=\text{softplus}\!\big(W^{(e)}_{\beta}P_{\beta}\big).$ (9)
A gating network $G(\mathbf{z}_{\text{WSI}})$ assigns softmax weights to experts, producing the time probability density(TPDF) and survival probability functions(SPF):
@@
The model is trained using negative log-likelihood loss handling both censored ($c=0$) and uncensored ($c=1$) data:
$\mathcal{L}_{\text{NLL}}=$ $-c\log\!\big(\text{TPDF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big)$ (12)
$-(1-c)\log\!\big(\text{SPF}(t_{d}\mid\mathbf{z}_{\text{WSI}})\big),$
combined with gating-entropy loss:
$\mathcal{L}_{\text{ent}}=-\sum_{e}G_{e}(\mathbf{z}_{\text{WSI}})\log(G_{e}(\mathbf{z}_{\text{WSI}}))$ (13)
to encourage diverse expert usage. The total loss is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{NLL}}+\lambda_{\text{ent}}\mathcal{L}_{\text{ent}}$.
@@ 4 Experiments and Results
We evaluate our framework on three TCGA cohorts: Lung Adenocarcinoma (LUAD; 459 WSIs), Kidney Renal Clear Cell Carcinoma (KIRC; 509 WSIs), and Breast Invasive Carcinoma (BRCA; 956 WSIs) [tcga2013pan]. Each dataset contains WSIs annotated with overall survival time, measured in months from diagnosis to death (uncensored) or last follow-up (censored). These cohorts represent diverse cancer types with distinct morphological characteristics, providing a comprehensive testbed for generalization.
@@
The quantile hyperparameter is set to $q=0.25$, retaining the top $75\%$ of patches based on logits. The $k$-NN graph constructed from the similarity matrix $S$ connects each point to its 10 nearest neighbors. Based on the graph connectivity, the patch indices are clustered into $G$ groups of 64 each, and these grouped patches are subsequently processed by an 8-head MHSA within the HCA module. The Expert-Driven Mixture of Log-Logistics module employs five experts, each modeling a log-logistic distribution (LLD) with $K=100$ components. The model is trained for 20 epochs with Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) and a batch size of 1. Results are reported as mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric over 5-fold cross-validation. All experiments were run on an NVIDIA A6000 GPU.
4.1 Comparative Performance Analysis
@@
Table 1a summarizes performance across TCGA datasets. Our framework consistently achieves the highest mean TDC, indicating superior discrimination and calibration. We compare against histology-based MIL models—AMIL [ilse2018attentionbaseddeepmultipleinstance], CLAM [lu2020dataefficientweaklysupervised], DSMIL [li2021dualstreammultipleinstancelearning], and TransMIL [shao2021transmiltransformerbasedcorrelated]—which served as backbones for the expert-driven mixture of log-logistics block. Survival-specific MIL methods such as HIPT [chen2022scalingvisiontransformersgigapixel], PatchGCN [chen2021slideimages2dpoint], HGT [10.1007/978-3-031-43987-2_72], SCMIL [Yang_2024] and OTSurv [RenQin_OTSurv_MICCAI2025] act as baselines. Since multimodal frameworks [10981028, ZhoHai_CoC_MICCAI2025, CuiJia_HiLa_MICCAI2025] lack public code, the results are taken from respective papers.
Our method attains TDC scores of $0.644\pm 0.059$ on LUAD, $0.751\pm 0.037$ on KIRC, and $0.752\pm 0.011$ on BRCA, achieving state-of-the-art performance on all datasets. The consistent improvement across diverse cancer types demonstrates the robustness of our approach. By combining quantile-based patch selection, graph-guided clustering, hierarchical attention, and expert-driven mixture of log-logistics, our model effectively identifies survival-relevant regions. Remarkably, it surpasses multimodal systems while relying solely on histopathology images.
4.2 Ablation Studies and Component Analysis